if a wire of length L is stretched from a fraction inorder to make its resistance four times the original
such that its new length is 3/2 times the original length...

then find the fraction by which it is increased....

Grade:12

1 Answers

mani kanta
11 years ago

Formula:

(R1/R2)*(l2/l1)=(A1/A2)

 

A1/A2=3/8